OpenStudy (anonymous):

what is the mean of 18 and 2? Thank you!

OpenStudy (anonymous):

The Mean is same to The average http://www.purplemath.com/modules/meanmode.htm Mean of 18 and 2 is (18+2)/2 = 10

OpenStudy (anonymous):

10 its the average 18+2=20 /2=10